OpenStudy (anonymous):

will fan and medal Which point lies on a circle that is centered at A(3, 3) and passes through B(6, 5)? C(1, 6) D(6, 0) E(0, 3) F(3, -1) G(3, 6)

OpenStudy (owlcoffee):

You'll have to find the equation of the circle and find a point belonging to it.

OpenStudy (anonymous):

first find the radius by finding the distance between the center and point on the circumference.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[r=\sqrt{\left( x _{2}-x _{1} \right)^2+\left( y _{2}-y _{1} \right)^2}\]

OpenStudy (anonymous):

here x1=3,y1=3 y1=6,y2=5

OpenStudy (anonymous):

ok so how do i get the answer?

OpenStudy (anonymous):

x2=6

OpenStudy (anonymous):

\[r=\sqrt{\left( 6-3 \right)^2+\left( 5-3 \right)^2}=?\]

OpenStudy (anonymous):

so wats

OpenStudy (anonymous):

\[r=\sqrt{9+4}=\sqrt{13}\]

OpenStudy (anonymous):

ok these are the possible answers which 1 is the answer C(1, 6) D(6, 0) E(0, 3) F(3, -1) G(3, 6

OpenStudy (owlcoffee):

We are not allowed to give out answers.

OpenStudy (anonymous):

eq. of circle is \[\left( x-h \right)^2+\left( y-k \right)^2=r^2\] where (h,k) is the center.

OpenStudy (anonymous):

ya you are just give me da answer i wont tell